OpenStudy (anonymous):

Simplify the following boolean algebraic equation ________ ________ Y=A.B'(A+C)+A'.B(A+B'+C')

OpenStudy (anonymous):

you can simplify it by using k-map.

OpenStudy (anonymous):

Y=A.B'(A+C)+A'.B(A+B'+C') =(AB'+AB'C)+(A'AB+A'BB'+A'BC') =AB'(1+C)+A'BC' =AB'+A'BC' =AB'+A'BC' Let V=AB' and W=A'BC' Y=V+W=((V+W)')' (Double Negation) Using DeMorgan's Theorems (M+N)'=M'.N' (M.N)'=(M+N)' Y=(V'.W' )' =((AB')'.(A'BC')')' =((A'+B).(A+B'+C))' =(A'A+A'B'+A'C+AB+BB'+BC)' =(A'B'+A'C+AB+BC)' =(P+Q+R+S)' =P'Q'R'S' =(A+B).(A+C').(A'+B).(B'+C') =(A+AB+AC').(A'B'+A'C'+BB'+BC') =(A+AC').(A'B'+A'C'+BC') =C'.(A'B'+A'C'+BC') =(A'B'C'+A'C')+BC' =A'C'(B'+1)+BC' =A'C'+BC' =C'(A'+B)

OpenStudy (anonymous):

Sorry I did a mistake. Y =(A+AC').(A'B'+A'C'+BC') =A.(A'B'+A'C'+BC') >>A Instead of C' =AA'B'+AA'C'+ABC' =ABC'